OpenStudy (anonymous):

Find the volume of a cube with the given side length. 8 units A. 24 cubic units B. 64 cubic units C. 512 cubic units D. 4096 cubic units

OpenStudy (anonymous):

@Study312

OpenStudy (anonymous):

@calculusxy

OpenStudy (study312):

8 x 8 x 8 = ?

OpenStudy (anonymous):

=512

OpenStudy (study312):

that's your answer. V=a3=83=512

OpenStudy (anonymous):

What is the volume of a cube with a side of length 11 units? A. 33 cubic units B. 121 cubic units C. 1331 cubic units D. 14,641 cubic units

OpenStudy (study312):

11 x 11 x 11 =?

OpenStudy (anonymous):

would i do11x11x11

OpenStudy (study312):

yes

OpenStudy (anonymous):

=1331

OpenStudy (study312):

yes. The formula is V = \[x ^{3}\]

OpenStudy (anonymous):

How many 1 by 1 unit cubes would equal the volume of the figure shown? A. 60 B. 120 C. 250 D. 3500

OpenStudy (calculusxy):

what is the figure?

OpenStudy (anonymous):

http://prntscr.com/9nbvwq

OpenStudy (study312):

Calculusxy why don't you take this one because I have no idea

OpenStudy (anonymous):

well he left

OpenStudy (calculusxy):

find the volume of cube one (large) \(volume = s^3\) add it with the volume of cube two (small) -- use the same formula of \(volume = s^3\)

OpenStudy (anonymous):

=3375

OpenStudy (study312):

I am going to do a history paper so I will be off and on, just thought you should know.

OpenStudy (anonymous):

=15

OpenStudy (anonymous):

ok when

OpenStudy (calculusxy):

the volume of the larger cube is 3375 cubic units what about the smaller one ?

OpenStudy (anonymous):

=125

OpenStudy (calculusxy):

so what's 3375 + 125?

OpenStudy (study312):

right now so If I am unresponsive give me a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

=3500

OpenStudy (calculusxy):

correct!

OpenStudy (anonymous):

What is the volume of the figure? A. 648 cubic units B. 216 cubic units C. 108 cubic units D. 54 cubic units

OpenStudy (anonymous):

http://prntscr.com/9nbyq9

OpenStudy (anonymous):

@calculusxy

OpenStudy (calculusxy):

notice how there is 3 congruent cube. can you see that?

OpenStudy (anonymous):

ummmmmm no

OpenStudy (anonymous):

@civicsiscool44

OpenStudy (civicsiscool44):

yes

OpenStudy (anonymous):

ok

OpenStudy (calculusxy):

now can you see it ?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

okay. since all three have the same side lengths (congruent) you can find the volume of one cube and then multiply it by 3

OpenStudy (civicsiscool44):

whoa holy moly thats confusing question

OpenStudy (anonymous):

lol

OpenStudy (calculusxy):

we know a cube has the same side lengths on all sides this case it is 6 \(\text{Volume of Cube} = s^3\) \(\text{Volume of Cube} = 6^3 = 6 \times 6 \times 6\)

OpenStudy (calculusxy):

did you get it ?

OpenStudy (anonymous):

=1296

OpenStudy (calculusxy):

not really. what is \(6^3\) ?

OpenStudy (anonymous):

216

OpenStudy (calculusxy):

good now what is \(216 \times 3\)?

OpenStudy (anonymous):

648

OpenStudy (anonymous):

is that right

OpenStudy (calculusxy):

yes !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

who plays call of duty

OpenStudy (study312):

I don't play call of duty. I play Kingdom Hearts and sometimes pokemon

OpenStudy (study312):

isn't 81 because volume of 1 = 3•3•3 = 27, 27 • 3 = 81?